The Jeter Corporation is considering acquiring the A-Rod Corporation. The data for the two companies are as follows:

A-Rod Corp. Jeter Corp.
Total earnings $ 708,000 $ 3,900,000
Number of shares of stock outstanding 295,000 1,950,000
Earnings per share $ 2.40 $ 2.00
Price-earnings ratio (P/E) 20 24
Market price per share $ 48 $ 48

a. The Jeter Corp. is going to give A-Rod Corp. a 50 percent premium over A-Rod’s current market value. What price will it pay? (Do not round intermediate calculations and round your answer to the nearest whole number.)
  

b. At the price computed in part a, what is the total market value of A-Rod Corp.? (Use the number of A-Rod Corp. shares times price.) (Do not round intermediate calculations and round your answer to the nearest whole number.)
  

c. At the price computed in part a, what is the P/E ratio Jeter Corp. is assigning A-Rod Corp? (Do not round intermediate calculations and round your answer to 2 decimal places.)
  

d. How many shares must Jeter Corp. issue to buy the A-Rod Corp. at the total value computed in part b? (Do not round intermediate calculations and round your answer to the nearest whole number.)
  

e. Given the answer to part d, how many shares will Jeter Corp. have after the merger? (Do not round intermediate calculations and round your answer to the nearest whole number.)
  

f. What are the new postmerger earnings per share? (Do not round intermediate calculations and round your answer to 2 decimal places.)
Answer #1
Solution:
a. Price   $72
working Notes:
Jeter Corp. is going to give A-Rod Corp. a 50 percent premium over A-Rod’s current market value means 50% above current price of A-Rod corp. which is $48
Hence Price it will pay to A-Rod Corp
= current market price of A-Rod   + 50 % x current market price of A-Rod
= $48   + 50 % x 48
= $48   + $24
=$72
b. Market value $21,240,000
Working Notes:
the total market value of A-Rod Corp. based on a. price
= Total no of shares outstanding of A-Rod corp. x price per share calculated in a.
= 295,000 x $72
=$21,240,000
c. P/E ratio   30 times
Working Notes:
P/E ratio assigned to A-Rod based on price calculated in a. above.
Price per share = $72
Earnings per share of A-Rod Corp is $2.40
P/E ratio = Price / Earnings price per share
P/E ratio = $72/ $2.40
P/E ratio = 30 times
d. New shares 442,500 shares
Working Notes:
New shares to be issued to A-Rod by Jeter Corp
= Total market value of A-Rod calculated in b. Above / Current market price of Jeter
=$21,240,000/$48
=442,500 shares
e. Total Shares 2,392,500    
Working Notes:
No of shares will Jeter Corp. have after the merger
= old no of shares outstanding in Jeter before merger + No of new shares to issued to A-Rod in d.
=1,950,000 + 442,500
=2,392,500
f. New post merger EPS    $1.93
Working notes:
New post merger EPS
= Total Earnings post merger / Total no of shares post merger
Total no of shares post merger = 2,392,500 calculated in e. above
Total Earnings post merger = Total earnings pre merger of both corp.
Total Earnings post merger = $708,000 + $3,900,000
Total Earnings post merger =$4,608,000
New post merger EPS
= Total Earnings post merger / Total no of shares post merger
=$4,608,000 /2,392,500
=$1.926018809
=$1.93
Please feel free to ask if anything about above solution in comment section of the question.